OpenStudy (anonymous):

what's the equation Keith can finish a job in 10 days and the same job would take Kenneth 6 days. How long would it take Karl working alone, if working together, all three can finish the job in 2 day?

OpenStudy (anonymous):

A) inverse of required-time = productivity (work efficiency) B) Sum the individual productivities, and choose as ur unknown the productivity of KArl C) when you find it - take the inverse- you got time-of-Karl

OpenStudy (anonymous):

? anyone there ?....

OpenStudy (anonymous):

1/6 + 1/10 + x = 1/2

Directrix (directrix):

@danielle02 The "work problem" explanation at the following link may be helpful: http://www.purplemath.com/modules/workprob.htm

OpenStudy (anonymous):

Keith can do 1/10 in 1 day Kenneth 1/6 in 1 day Karl ? in 1 day Total 1/2 in 1day

OpenStudy (anonymous):

x = 14/60 = 7/30

OpenStudy (anonymous):

KArl time = 4 and 2/7 days

Directrix (directrix):

I'm thinking: 1/6 + 1/10 + 1/x = 1/2. x = 30/7 days --> 4 2/7 days as reported above by @Mikael

OpenStudy (anonymous):

tnx @danielle02

OpenStudy (anonymous):

thanks! @Mikael